Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Question from the 'GMAT Review 12th ed'

Expert replies
by bmaree » Tue Dec 06, 2011 2:54 am
I hope this is the right place to post this question.
I'm currently studying the 'Math Review' section of the Official Guide for GMAT Reivew. Thing were going pretty fine till I reached the last paragraph of the 'Discrete Probability' section on page 120.
For those of you who have the book, I would greatly appreciate it if someone could give me insight into how did they reach the conclusions in this paragraph.
I'll retype the question for easier reference.
Consider an experiment with events A, B and C for which P(A)= 0.23, P(B)= 0.40 and P(C)= 0.85,
Also suppose that A and B are mutually exclusive and B and C are independent.
P(A or B)= P(A)+P(B) since mutually exclusive
= 0.23+0.40 = 0.63
P(B or C)= P(B)+P(C)-P(B)P(C) by independence
=0.40+0.85 = 0.91
Note that P(A or C) and P(A and C) cannot be determined using the information given
[I assume that's because the question didn't state whether they're mutually exclusive or independent in relevance to each other]
But it can be determined that A and C are not mutually exclusive since P(A)+P(C)=1.08, which is greater than 1, and therefore cannot equal P(A or C); from this it follows that P(A and C)>= 0.08. [How did they reach this conclusion? where did the 0.08 come from?]
The rest of the paragraph is clear.
Thank you in advance for any help.
Join the discussion
Source: — Problem Solving |

by Ian Stewart » Tue Dec 06, 2011 11:10 am
bmaree wrote:
P(B or C)= P(B)+P(C)-P(B)P(C) by independence
=0.40+0.85 = 0.91
That last line should read:

=0.40+0.85 - (0.40)(0.85) = 0.91
bmaree wrote:
Note that P(A or C) and P(A and C) cannot be determined using the information given[/color] [I assume that's because the question didn't state whether they're mutually exclusive or independent in relevance to each other]
But it can be determined that A and C are not mutually exclusive since P(A)+P(C)=1.08, which is greater than 1, and therefore cannot equal P(A or C); from this it follows that P(A and C)>= 0.08. [How did they reach this conclusion? where did the 0.08 come from?]
If A and C were mutually exclusive, then P(A or C) would be equal to P(A) + P(C), so would be equal to 0.23 + 0.85 = 1.08. But that's greater than 1, and probabilities can never be greater than 1. So, since we get a nonsensical answer if we assume A and C are mutually exclusive, it must be impossible that A and C are mutually exclusive. Finally the 0.08 comes from the following: P(A or C) = P(A) + P(C) - P(A and C). We know that P(A or C) must be at most 1. Since P(A) + P(C) is 1.08, then to bring the right side of that equation down to a value of at most 1, it must be that P(A and C) is at least 0.08.

All of that said, I'd strongly recommend *not* using the OG to review counting and probability principles. Their treatment of the subject is pedantically formal and abstract. It's neither easy to follow, nor easy to apply to actual test questions. In general, the practice questions in the OG are great, but the review material at the beginning leaves a lot to be desired.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion